OpenStudy (anonymous):

Please help!!! Evaluate the function for the given values: f(x)=-2^3-4x^2 +5 at x=-3

OpenStudy (anonymous):

Well, we see that \[f(x)=(-2)^3-4x^2+5| x=-3\] So, \[f(-3)=-2^3-4*(-3)^2+5\] Further simplifying gives us \[f(-3)=-8-36+5=-39\]